OpenStudy (anonymous):

Part 1: Find the solution to the equation 64(3 – x) = 42x. (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (akashdeepdeb):

All you have to do in this question is open the brackets like this: 192 - 64x = 42x And then bring all x values to one side and then find x. That's pretty much it! :)

OpenStudy (akashdeepdeb):

Try using this concept in All. :)

OpenStudy (anonymous):

Opps now that I look at it, the problem didn't go the way it was suppose to

OpenStudy (anonymous):

64^(3-x) = 4^2x

OpenStudy (akashdeepdeb):

haha... Write 64 as 4^3 and multiply 3 to 64's power What i mean is that. 64^(3-x) can also be written like 4^3(3-x) And this is equal to 4^2 Now we know if BASES (here, 4) are equal then the powers should also be the SAME. So here 3(3-x) = 2 Now just solve it by openning the brackets and find the x balue! :) Got it? :)

OpenStudy (anonymous):

Thanks